OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)? y − 5 = 3(x − 1) y − 2 = −3(x − 0) y + 2 = 3(x + 1) y + 5 = −3(x + 1)

OpenStudy (anonymous):

The slope is 3, and you can use either set of points to plug into the equation y-y_1=m(x-x_1) so if we use 0 and 2 it would be y-2=3(x-0)

OpenStudy (anonymous):

So the first option is correct

OpenStudy (anonymous):

slope = m\[m=\frac{y-y_o}{x-x_o}\]\[m=\frac{2-5}{0-1}=3\]\[y-y_o=m(x-x_o)\]The answer is:\[y-5=3*(x-1)\]

OpenStudy (anonymous):

ok thank you guys so much